OpenStudy (anonymous):

We are given an input current Iin=100μA and a linear resistive load represented by the resistor RL . Our resistive load is not regulated for the high current provided by our current source, so we add a resistor RS in parallel to divide the current such that IL≈40μA. An additional requirement is that the Thevenin resistance as seen from the load terminals is between 60kΩ and 80kΩ. Assume first that the resistors have their nominal resistance. Come up with resistors RS and RL such that the divider ratio IL/Iin is within 10% of the requirement. Of course, the resistances you chose are just nomin

OpenStudy (anonymous):

I = E/R = (E/Rl) + (E/Rs) = E*[(1/Rl)+(1/Rs)] = 100 mA (E/Rl) = 40 mA (E/Rs) = 60 mA E = 40Rl = 60Rs Rl = (60/40)Rs = 1.5*Rs Take middle of 60 Ohms and 80 Ohms => 70 Ohms 70 = (Rl*Rs)/(Rl+Rs) = (1.5*Rs)*(Rs) / (1.5Rs + Rs) = (1.5*Rs) / 2.5 Rs = 70 (2.5/1.5) = 116.7 Ohms Rl = 1.5*Rs = 175 Ohms Resistor values have denominations of 10 12 15 18 22 27 33 39 47 56 68 and 82 Ohms Choose Rl = 182 Ohms Rs = 120 Ohms Check (Rl*Rs)/(Rl+Rs) = (180*120) / (180+120) = 72.3 Ohms